Circular No. 8870

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 R4 (GARRADD) G. J. Garradd reports his discovery of a comet on CCD images taken

with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey (discovery observation tabulated below), the object indicated as having a faint 20′′ tail to the west on 30-s exposures. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, A. C. Gilmore and P. M. Kilmartin (Mt. John 1.0-m f/7.7 reflector) write that their images from Sept. 16.50–16.56 UT show a coma of diameter ∼ 5′′ and a ‘wedge’-shaped tail 25′′ long in p.a. 270^{o}; Gilmore’s images taken on Sept. 17.5 show a coma diameter of ∼ 9′′ with central condensation (FWHM 4′′) and a tail 20′′ long in p.a. 260^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Sept.14.60795  2^{h}15^{m}02.^{s}73  −26^{o}34′54′′.8  18.4

The available astrometry (including prediscovery Siding Spring images from Aug. 12), together with the following preliminary elliptical orbital elements and an ephemeris, appear on MPEC 2007-S04.

T = 2007 Sept.26.904 TT  ω = 282.^{o}601
e = 0.67380  Ω = 87.603 }2000.0
q = 1.92336 AU  i = 20.240
a = 5.89621 AU n^{o} = 0.068841 P = 14.3 years

VARIABLE STAR IN PERSEUS H. Mikuž, Crni Vrh Observatory, reports that a new starlike variable

(mag 17.9, roughly equivalent to the R band) was detected on four unfiltered CCD images (limiting mag 19.0) taken around Sept. 16.078 UT in the course of the PIKA project at the Crni Vrh Observatory using the 0.60-m f/3.3 Cichocki telescope. The variable object is located at α = 3^{h}57^{m}28^{s}.01, δ = +35^{o}37′16′′.3 (equinox 2000.0). This position is consistent with USNO-B1-catalogue star 1256-0061977 with B magnitude 21.5. Additional PIKA magnitudes for the new object show that it is brightening: 2003 Sept. 26, [19.0; 2004 Feb. 9, [19.0; 2005 Jan. 11, [19.0; 2007 Sept. 16.955, B = 17.90 ± 0.10, R = 17.65 ± 0.05; Sept. 17.037, B = 17.60 ± 0.09, R = 17.39 ± 0.04. Nothing is visible at this location on several Palomar Sky Survey and 2MASS images from the Aladin database.

2007 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green